Equation 

x  y p  x  y q  z  z

2

cos( ) sin( )

is a quasi-linear equation


1
Expert's answer
2021-06-01T10:34:20-0400

Quasi-linear partial differential equations of first order can be written as


"P(x,y,z)p+Q(x,y,z)q=R(x,y,z)\quad\quad (1)"

where "P,Q,R" are functions of "x,y" and "z" . They do not involve "p" or "q" . Here, linear means that "p" and "q" appear to the first degree only.

In our case, the equation


"\cos(x+y)p+\sin(x+y)q=z^2"


is a quasi-linear equation.

 It's linear with respect to the derivatives (p and q)